PLEASE HELP
Find the effective rate corresponding to the following nominal rate of 6% compounded monthly.
Round to the nearest hundredth.

Answer by Alan3354(69443)

Find the effective rate corresponding to the following nominal rate of 6% compounded monthly.
Round to the nearest hundredth.
================
If you mean 6% per year:
----
(1 + 0.06/12)^12 = 1.005^12 = 1.0617
--> 6.17%